# Implementation of Leaking Quantum Walks on a Photonic Processor

### v1

@@ 3.2 Measurements
The quantum walker was injected at different initial modes within the lattice, either near the leaking boundary (waveguides 2 and 3) or near the reflecting boundary (waveguides 6 and 7). Two distinct leakage regimes were explored by tuning the reflection parameter to for weak leaking and for strong leaking. For each configuration, the walker was allowed to evolve for a number of steps ranging from 4 to 20. At each step, we measured the spatial probability distribution of the surviving wavefunction within the lattice. From these distributions, the mean position and the variance of the walker according to equations 2.1.
@@
The mean position and variance for the strong-leakage regime are shown in Figure 4. We observe good agreement between the experimental data (symbols) and the numerically simulated behavior (lines). The mean position reveals a net propagation of the walker across the lattice from its initial location toward the opposite boundary. When the walker is initialized near the leaking edge (input 2), it rapidly propagates toward the reflecting boundary, reaching an average displacement of approximately four lattice sites from the initial position after 15 steps (blue symbols in Figure 4, left panel). By contrast, when the walker is injected near the reflecting boundary (input 7), its propagation is initially slower; nevertheless, it attains a larger overall displacement, exceeding six lattice sites and reaching the position farthest from the initial mode after 16 steps.
The variance of the position reveals distinct spreading behaviors depending on the input site and the measurements (symbols) show a good agreement with the simulated curves (lines). However, here, when the walker is injected close to the reflecting boundary, the wavefunction exhibits limited spreading, as evidenced by the nearly flat variance curve for input 7 (cyan diamonds in Figure 4, right panel). For more central input sites, the wavefunction initially spreads and subsequently relocalizes (input 6, violet triangles), and in some cases spreads again (orange squares). When the input position is close to the leaking boundary, the wavefunction still displays spreading and relocalization dynamics; however, the amplitude of the spreading is reduced (blue dots).
@@
We also measured the same quantities (mean position of the walker and its variance) in the case of a weak leaking boundary, by setting the parameters of our photonic processor such that the reflectivity of the beam splitters at the boundary is . We measured the output probability distribution of the walker when injected in waveguide 2 and 6, close to the leaking and the reflecting boundaries respectively. Results are reported in Figure 5. Here we can observe a good agreement between experimental (symbols) and simulated (lines) data. It is interesting to note that if the walker initial position is close to the absorbing boundary, it moves toward the other edge more slowly than in the case of strong leaking, reaching the farthest position of its walk at step 19 (blue circles in Figure 5 left panel). However its displacement is larger, reaching a value close to (compared to in the previous configuration). For initial position corresponding to waveguide 6, the walker’s mean position oscillates with a small displacement of around (violet triangles). When looking at the variance of the mean position, a similar behavior to the configuration of strong leaking is observed: if the initial position is close to the reflecting boundary the wavefunction spreads and relocalizes leading to an oscillating behavior of the variance (violet triangles in Figure 5 right panel), while a limited spreading almost without oscillations if the walker enters the lattice close to the leaking boundary (blue circles).
As a final comparison, we analyze the mean position and variance of the walker wavefunction for two different input sites, namely waveguides 2 and 6, in two different configurations of leakage probability. These behaviors are compared with numerical simulations of a quantum walk on a lattice without leakage, corresponding to fully reflective boundary conditions at both edges. The results are shown in Figures 6 and 7 for inputs 2 and 6, respectively.
@@
As predicted by the simulations discussed in the previous section, when the input site is close to the leaking edge, the weak-leakage regime (cyan triangles and dashed line in Figure 6, left panel) closely reproduces the behavior observed in the absence of leakage (orange dotted line). In contrast, at higher leakage probabilities (blue dots and solid blue line), the dynamics differs significantly: the walker propagates more rapidly and reaches its maximum displacement at approximately step 13, although this displacement is reduced with respect to the other two cases compared to . This behavior is also reflected in the variance (Figure 6, right panel), where the strong-leakage regime exhibits oscillations with a larger amplitude (blue circles and solid line) than those observed in the no-leakage or weak-leakage cases (orange dotted line and cyan triangles with dashed line, respectively).
@@
When the input site is far from the leaking boundary, a markedly different behavior is observed (Figure 7). During the first eight steps of the walk, the dynamics are nearly identical across all three regimes. Subsequently, in the strong-leakage case (violet diamonds and solid line), the mean position exhibits oscillations with a higher frequency compared to the other two regimes, whose behavior remains almost unchanged (Figure 7, left panel). By contrast, the variance displays very similar dynamics for all three regimes up to 18 steps (Figure 7, right panel), with only a slightly faster relocalization observed in the strong-leakage case between steps 8 and 13.
@@ 4 Discussion
The comparison between different leaking regimes for the injection mode 2, which is the closest to the lossy boundary, highlights the strongest impact of these effects. While weak losses leave the dynamics largely unchanged with respect to the lossless case, strong leakage leads to faster propagation of the mean position as well as enhanced spreading of the wavefunction. Additional features of the long-time dynamics, including the emergence of time windows characterized by enhanced or reduced variance and flattening of the mean displacement, are discussed in Appendix B.
Dynamical differences between the two limiting scenarios, i.e. absence of losses and a fully absorbing boundary, are particularly pronounced for the injection site closest to the leaking edge, conforming the key role played by boundary proximity in shaping both the early- and long-time dynamics. A detailed comparison between the lossless and fully absorbing boundary conditions is reported in Appendix B.
@@
Overall, these findings are confirmed by experimental measurements performed on the photonic processor, which show good agreement with the simulated mean position and variance across different input sites and leakage regimes (Figures 4, 5, 6, 7). Importantly, the experiments allow a closer inspection of the first 20 steps, highlighting subtle differences in the early-time propagation that are less visible in the long-time simulations.
@@
Finally, it is worth noting that, in both the numerical model and its experimental implementation, we assumed a constant probability for a photon reaching the leaking boundary to escape the lattice at each time step. This choice allowed us to isolate the effect of spatially localized losses, avoiding to superimpose contributions from their time dependence. If the loss coefficients (i.e., the BS reflectivity at the boundary) were step-dependent, different propagation paths would be unevenly weighted, likely resulting in a partial suppression of the interference effects and potentially introducing temporal memory in the interference pattern. From an experimental perspective, assuming the leakage probability independent on the specific time step at which the walker reaches the boundary corresponds to an idealized scenario which was expected to be more controllable and suitable for a first investigation of the role played by a localized partially absorbing boundary. A more general and systematic study of the effects of time-dependent or randomly fluctuating losses, which would model more complex and realistic scenarios, is left for future works.
@@ 5 Conclusions
FundingThis research was funded by QuantERA II Programme supported by the EU H2020 research and innovation Programme under GA No 101017733, with funding Italian organization PNRR MUR project PE0000023-NQSTI (Spoke 6, CUP: H43C22000870001).
@@
AcknowledgmentsThe authors acknowledge Caterina Taballione from Quix Quantum, Marco Barbieri and Ilaria Gianani from Roma Tre University for fruitful and constructive conversations.
Conflicts of interestThe authors declare no conflicts of interest.
@@ Appendix A Long-time numerical simulations
This appendix provides an extended analysis of the numerical simulations for the DTQW on a finite lattice in the presence of a leaking boundary, focusing on the long-time dynamics up to time steps. In particular, we report here on those dynamical features that were only briefly mentioned in Section 4, with the aim of supporting and clarify our physical interpretation. These results complement the main text by providing a detailed comparison between different leakage regimes and injection sites.
As reported in Section 4, the effects of losses strongly depend on the position of the injection site, as well as on the leakage strength. Figure 8 therefore summarizes the time evolution of the walker’s mean position and variance over steps for different injection sites and for the two representative leakage probabilities, namely (weak leakage) and (strong leakage). For reference, the lossless case () and the fully absorbing boundary () are also shown.
We first focus on the injection site closest to the leaking boundary (input 2), which in the main text was identified as the configuration most sensitive to losses. In the weak-leakage regime, the dynamics closely resemble those of the lossless case over several tens of steps, with only moderate differences in amplitude. By contrast, strong leakage induces significant deviations already at early times, including a faster propagation of the mean position and an enhanced spreading of the wavefunction, as evidenced by the increased variance. For injection sites progressively farther from the leaking boundary (input sites 3, 6, and 7), the impact of losses is progressively reduced. In these configurations, both the mean position and the variance exhibit similar temporal trends across different leakage regimes over a large number of steps, indicating that the walker dynamics remain largely governed by coherent propagation within the lattice.
To further clarify the role of the leaking regimes discussed above, it is instructive to analyze the two limiting cases of fully reflective and fully absorbing boundaries. The lossless case represents an ideal quantum walk confined between two perfectly reflecting boundaries. In this regime, the mean position exhibits large-amplitude oscillations, and the variance displays sharp peaks associated with strong interference effects and transient refocusing of the wavefunction. In the opposite limit of a fully absorbing boundary, most walkers are lost before contributing significantly to the spatial spreading. As a result, the variance remains low for an extended time window, with pronounced peaks appearing only at later times due to the residual fraction of the surviving wavefunction reaching the lattice edges. These limiting cases provide useful benchmarks for interpreting the intermediate leakage regimes discussed in the main text.
@@ Appendix B Extended discussion on long-time dynamics
Let’s start our discussion from the injection site 2. Interestingly, if the walker is initialized closed to the leaking boundary, time windows emerge in which the mean position exhibits a less steep trend and stays close to the center of the lattice, while the variance reaches large values (e.g., between time steps 55 and 75, where the red curve in Figure 8 (a) seems to lack a peak, while the green curve in Figure 8 (b) displays its highest peak where the orange curve has a drop). This observation indicates that, whereas the absence of losses or their weak presence would lead to photon localization or bunching, strong leakage can suppress localization effects inducing a broader distribution of the probability amplitude across the waveguides while keeping the average position close to zero.
@@
Injection site 2, due to its proximity to the leaking edge, results in the most pronounced dynamical differences between the two limiting scenarios, namely the absence of losses () and the presence of a fully absorbing boundary (), with a leaking probability equal to 1. The lossless case is the ideal QW, where no photons escape the system: the walker’s mean position exhibits the widest oscillations (dashed line in Figure 8 (a)), indicating that the wavefunction is reflected between the lattice edges and interference preserves a high degree of coherence. This remark is reinforced by variance (dashed line in Figure 8 (b)), that displays sharp peaks reaching values above 8. This implies that photons do not merely diffuse, but rather interfere, giving rise to pronounced interference lobes in the probability distribution. The sudden drops in variance point to transient re-focusing events, in which the wavefunction becomes more localized before spreading again. By contrast, in the total-loss regime the variance remains low for an extended time window (up to about the first 55 steps), indicating that most walkers are lost before contributing significantly to the spatial spreading. The appearance of a sharp peak in the variance at later times (beyond approximately 75 steps) can be attributed to the residual fraction of the surviving wavefunction reaching the lattice edges, which results in an increase in the computed variance even though the overall signal intensity is strongly reduced.
@@
For the injection site maximally distant from the leaking boundary, i.e. input site 7, a counterintuitive behavior emerges at long time: although far from the absorbing edge leakage would be expected to have minimal influence over long times, weak losses cause a flattening of the mean position (between steps 65 and 90), while stronger losses preserve the oscillatory pattern. This effect can be ascribed to the fact that weak losses do not strongly suppress wavefunctions reaching the leaking edge, but instead partially attenuate and reinject them into the dynamics. This process breaks the interference symmetry and induces an effective directional bias. The corresponding variance dynamics supports this interpretation. After an initial localization phase common to all regimes during the fist 20-25 steps, spreading occurs with reduce amplitude and broader peaks compared to those observed for more internal injection sites. Moreover, at longer times, variance of the high-leakage case presents a marked drop around the 50th time step, revealing a transient relocalization of the wavefunction that is absent when losses are weaker. This behavior mirrors that observed for the injection site 2, albeit with inverted roles between low- and high-loss cases, and underscores the nontrivial interplay between distance from the leaking boundary and loss-induced interference effects.

### v2

@@ 3.2 Measurements
The quantum walker was injected at different initial modes within the lattice, either near the leaking boundary (waveguides 2 and 3) or near the reflecting boundary (waveguides 6 and 7). Two distinct leakage regimes were explored by tuning the reflection parameter to for weak leaking and for strong leaking. For each configuration, the walker was allowed to evolve for a number of steps ranging from 4 to 20. At each step, we measured the spatial probability distribution of the surviving wavefunction within the lattice. From these distributions, the mean position and the variance of the walker according to Equations 4.
@@
The mean position and variance for the strong-leakage regime are shown in Figure 5. We observe good agreement between the experimental data (symbols) and the numerically simulated behavior (lines). The mean position reveals a net propagation of the walker across the lattice from its initial location toward the opposite boundary. When the walker is initialized near the leaking edge (input 2), it rapidly propagates toward the reflecting boundary, reaching an average displacement of approximately four lattice sites from the initial position after 15 steps (blue circles in Figure 5, left panel). By contrast, when the walker is injected near the reflecting boundary (input 7), its propagation is initially slower; nevertheless, it attains a larger overall displacement, exceeding six lattice sites and reaching the position farthest from the initial mode after 16 steps (green diamonds). The variance of the position reveals distinct spreading behaviors depending on the input site and the measurements (symbols) show a good agreement with the simulated curves (lines). However, here, when the walker is injected close to the reflecting boundary, the wavefunction exhibits limited spreading, as evidenced by the nearly flat variance curve for input 7 (green diamonds in Figure 5, right panel). For more central input sites, the wavefunction initially spreads and subsequently relocalizes (input 6, violet triangles), and in some cases spreads again (orange squares). When the input position is close to the leaking boundary, the wavefunction still displays spreading and relocalization dynamics; however, the amplitude of the spreading is reduced (blue circles).
@@
We also measured the same quantities (mean position of the walker and its variance) in the case of a weak leaking boundary, by setting the parameters of our photonic processor such that the transmissivity of the beam splitters at the boundary is . We measured the output probability distribution of the walker when injected in waveguide 2 and 6, close to the leaking and the reflecting boundaries respectively. Results are reported in Figure 6. Here we can observe a good agreement between experimental (symbols) and simulated (lines) data. It is interesting to note that if the walker initial position is close to the absorbing boundary, it moves toward the other edge more slowly than in the case of strong leaking, reaching the farthest position of its walk at step 19 (cyan circles in Figure 5 left panel). However its displacement is larger, reaching a value close to (compared to in the previous configuration). For initial position corresponding to waveguide 6, the walker’s mean position oscillates with a small displacement of around (pink triangles). When looking at the variance of the mean position, a similar behavior to the configuration of strong leaking is observed: if the initial position is close to the reflecting boundary, the wavefunction spreads and relocalizes leading to an oscillating behavior of the variance (pink triangles in Figure 6 right panel), while a limited spreading almost without oscillations is observed if the walker enters the lattice close to the leaking boundary (cyan circles).
@@
As a final comparison, we analyze the mean position and variance of the walker wavefunction for two different input sites, namely waveguides 2 and 6, in two different configurations of leakage probability. These behaviors are compared with numerical simulations of a quantum walk on a lattice without leakage, corresponding to fully reflective boundary conditions at both edges. The results are shown in Figures 7 and 8 for inputs 2 and 6, respectively. As predicted by the simulations discussed in the previous section, when the input site is close to the leaking edge, the weak-leakage regime (cyan triangles and solid line in Figure 7, left panel) closely reproduces the behavior observed in the absence of leakage (gray dash-dotted line). In contrast, at higher leakage probabilities (blue dots and dashed blue line), the dynamics differs significantly: the walker propagates more rapidly and reaches its maximum displacement at approximately step 13, although this displacement is reduced with respect to the other two cases compared to . This behavior is also reflected in the variance (Figure 7, right panel), where the strong-leakage regime exhibits oscillations with a larger amplitude (blue circles and dashed line) than those observed in the no-leakage or weak-leakage cases (gray dash-dotted line and cyan triangles with solid line, respectively).
@@
When the input site is far from the leaking boundary, a markedly different behavior is observed (Figure 8). During the first 8 steps of the walk, the dynamics are nearly identical across all three regimes. Subsequently, in the strong-leakage case (violet circles and dashed line), the mean position exhibits oscillations with a higher frequency compared to the other two regimes, whose behavior remains almost unchanged (Figure 8, left panel). By contrast, the variance displays very similar dynamics for all three regimes up to 18 steps (Figure 8, right panel), with only a slightly faster relocalization observed in the strong-leakage case between steps 8 and 13.
@@ 4 Discussion
The comparison between different leaking regimes for the injection mode 2, which is the closest to the lossy boundary, highlights the strongest impact of these effects. While weak losses leave the dynamics largely unchanged with respect to the lossless case, strong leakage leads to faster propagation of the mean position as well as enhanced spreading of the wavefunction. Additional features of the long-time dynamics, including the emergence of time windows characterized by enhanced or reduced variance and flattening of the mean displacement, are discussed in Appendix B. Dynamical differences between the two limiting scenarios, i.e. absence of losses and a fully absorbing boundary, are particularly pronounced for the injection site closest to the leaking edge, confirming the key role played by boundary proximity in shaping both the early- and long-time dynamics. A detailed comparison between the lossless and fully absorbing boundary conditions is also reported in Appendix B.
@@
Overall, these findings are confirmed by experimental measurements performed on the photonic processor, which show good agreement with the simulated mean position and variance across different input sites and leakage regimes (Figures 5, 6, 7, 8). Importantly, the experiments allow a closer inspection of the first 20 steps, highlighting subtle differences in the early-time propagation that are less visible in the long-time simulations.
@@
Finally, it is worth noting that, in both the numerical model and its experimental implementation, we assumed a constant probability for a photon reaching the leaking boundary to escape the lattice at each time step. This choice allowed us to isolate the effect of spatially localized losses, avoiding to superimpose contributions from their time dependence. If the loss coefficients (i.e., the BS transmissivity at the boundary) were step-dependent, different propagation paths would be unevenly weighted, likely resulting in a partial suppression of the interference effects and potentially introducing temporal memory in the interference pattern. From an experimental perspective, assuming the leakage probability independent of the specific time step at which the walker reaches the boundary corresponds to an idealized scenario which was expected to be more controllable and suitable for a first investigation of the role played by a localized partially absorbing boundary. A more general and systematic study of the effects of time-dependent or randomly fluctuating losses, which would model more complex and realistic scenarios, is left for future works.
@@ 5 Conclusions
Funding: This research was funded by QuantERA II Programme supported by the EU H2020 research and innovation Programme under GA No 101017733, with funding Italian organization PNRR MUR project PE0000023-NQSTI (Spoke 6, CUP: H43C22000870001).
@@
Acknowledgments: The authors acknowledge Caterina Taballione from Quix Quantum, Marco Barbieri and Ilaria Gianani from Roma Tre University for fruitful and constructive conversations.
Conflicts of interest: The authors declare no conflicts of interest.
@@ Appendix A Long-time numerical simulations
This appendix provides an extended analysis of the numerical simulations for the DTQW on a finite lattice in the presence of a leaking boundary, focusing on the long-time dynamics up to time steps. In particular, we report here on those dynamical features that were only briefly mentioned in Section 3.1, with the aim of supporting and clarify our physical interpretation. These results complement the main text by providing a detailed comparison between different leakage regimes and injection sites.
As reported in Section 3.1, the effects of losses strongly depend on the position of the injection site, as well as on the leakage strength. Figure 9 therefore summarizes the time evolution of the walker’s mean position and variance over steps for different injection sites and for the two representative leakage probabilities, namely (weak leakage) and (strong leakage). For reference, the lossless case () and the fully absorbing boundary () are also shown.
We first focus on the injection site closest to the leaking boundary (input 2, shown in Figure 9 (a)-(b)), which in the main text was identified as the configuration most sensitive to losses. In the weak-leakage regime (cyan solid line), the dynamics closely resemble those of the lossless case (gray dotted line) over several tens of steps, with only moderate differences in amplitude. By contrast, strong leakage (blue dashed line) induces significant deviations already at early times, including a faster propagation of the mean position and an enhanced spreading of the wavefunction, as evidenced by the increased variance. For injection sites progressively farther from the leaking boundary (input sites 3, 6, and 7 shown in Figure 9 (c)-(d), (e)-(f), (g)-(h), respectively), the impact of losses is progressively reduced. In these configurations, both the mean position and the variance exhibit similar temporal trends across different leakage regimes over a large number of steps, indicating that the walker dynamics remain largely governed by coherent propagation within the lattice.
To further clarify the role of the leaking regimes discussed above, it is instructive to analyze the two limiting cases of fully reflective (transmissivity ) and fully absorbing (transmissivity ) boundaries on input 2, represented in Figure 9 as gray dotted and black dashed-dotted lines, respectively. These limiting cases provide useful benchmarks for interpreting the intermediate leakage regimes discussed in the main text.
@@ Appendix B Extended discussion on long-time dynamics
Let’s start our discussion from the injection site 2. Interestingly, if the walker is initialized closed to the leaking boundary, time windows emerge in which the mean position exhibits a less steep trend and stays close to the center of the lattice, while the variance reaches large values (e.g., between time steps 55 and 75, where the blue dashed curve in Figure 9 (a) seems to lack a peak, while the blue dashed curve in Figure 9 (b) displays its highest peak where the cyan solid line has a drop). This observation indicates that, whereas the absence of losses or their weak presence would lead to photon localization or bunching, strong leakage can suppress localization effects inducing a broader distribution of the probability amplitude across the waveguides while keeping the average position close to zero. Injection site 2, due to its proximity to the leaking edge, results in the most pronounced dynamical differences between the two limiting scenarios, namely the absence of losses (transmissivity ) and the presence of a fully absorbing boundary (transmissivity ), with a leaking probability equal to 1. The lossless case is the ideal QW, where no photons escape the system: the walker’s mean position exhibits the widest oscillations (gray dotted line in Figure 9 (a)), indicating that the wavefunction is reflected between the lattice edges and interference preserves a high degree of coherence. This remark is reinforced by variance (gray dotted line in Figure 9 (b)), that displays sharp peaks reaching values above 8. This implies that photons do not merely diffuse, but rather interfere, giving rise to pronounced interference lobes in the probability distribution. The sudden drops in variance point to transient re-focusing events, in which the wavefunction becomes more localized before spreading again. By contrast, in the total-loss regime (black dash-dotted line) the variance remains low for an extended time window (up to about the first 55 steps), indicating that most walkers are lost before contributing significantly to the spatial spreading. The appearance of a sharp peak in the variance at later times (beyond approximately 75 steps) can be attributed to the residual fraction of the surviving wavefunction reaching the lattice edges, which results in an increase in the computed variance even though the overall signal intensity is strongly reduced.
@@
For the injection site maximally distant from the leaking boundary, i.e. input site 7, a counterintuitive behavior emerges at long time: although far from the absorbing edge leakage would be expected to have minimal influence over long times, weak losses cause a flattening of the mean position (between steps 65 and 90), as displayed by the green solid line in Figure 9 (g), while stronger losses (green dashed line) preserve the oscillatory pattern. This effect can be ascribed to the fact that weak losses do not strongly suppress wavefunctions reaching the leaking edge, but instead partially attenuate and reinject them into the dynamics. This process breaks the interference symmetry and induces an effective directional bias. The corresponding variance dynamics supports this interpretation (Figure 9 (h)). After an initial localization phase common to all regimes during the fist 20-25 steps, spreading occurs with reduce amplitude and broader peaks compared to those observed for more internal injection sites. Moreover, at longer times, variance of the high-leakage case (green dashed line) presents a marked drop around the 50th time step, revealing a transient relocalization of the wavefunction that is absent when losses are weaker (green solid line). This behavior mirrors that observed for the injection site 2, albeit with inverted roles between low- and high-loss cases, and underscores the nontrivial interplay between distance from the leaking boundary and loss-induced interference effects.